Cambridge University Cambridge University Cambridge University Cambridge University

First Printed: November 1989 (for HOL88 Version 1.11)

Revised Edition: July 1991 (for HOL88 Version 2) Revised Edition: March 1994 (for HOL88 Version 2.02)

First Printed: November 1989 (for HOL88 Version 1.11)

Revised Edition: July 1991 (for HOL88 Version 2) Revised Edition: March 1994 (for HOL88 Version 2.02)

First Printed: November 1989 (for HOL88 Version 1.11)

Revised Edition: July 1991 (for HOL88 Version 2) Revised Edition: March 1994 (for HOL88 Version 2.02)

First Printed: November 1989 (for HOL88 Version 1.11)

Revised Edition: July 1991 (for HOL88 Version 2) Revised Edition: March 1994 (for HOL88 Version 2.02)